OpenStudy (anonymous):

When the sum of a number plus 3 is squared, it is 11 more than the sum of the number plus 2 when squared.

Parth (parthkohli):

"a number" ---> x \( \color{Black}{\Rightarrow (x + 3)^2 = 11 + (x + 2)^2}\)

OpenStudy (anonymous):

What is the number though?

Parth (parthkohli):

You get the number when you solve the equation.

OpenStudy (anonymous):

How would i solve it.. sorry but i suck at math (hints the name)

Parth (parthkohli):

\( \color{Black}{\Rightarrow x^2 + 6x + 9 = 11 + x^2 + 4x + 4}\)

Parth (parthkohli):

\( \color{Black}{\Rightarrow x^2 + 6x + 9 = 15 + x^2 + 4x}\) \( \color{Black}{\Rightarrow 2x - 6 = 0}\)

Parth (parthkohli):

\( \color{Black}{\Rightarrow 2x = 6}\) \( \color{Black}{\Rightarrow x = 6 \div 2}\)

OpenStudy (anonymous):

Thanks